OpenStudy (anonymous):

Use the Substitution Method to solve the following system of equations. 4x + y = 11 x + 2y = 8 (3, 2) (2, 3) (-2, -3) (-2, 3)

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

4x+y=11 x+2y=8 take the second equation: x+2y=8 make it x= so x=8-2y then plug that back into the first equation 4(8-2y)+y=11 32-8y+y=11 32-7y=11 -7y=-21 7y=21 y=3 now plug y into the second equation and solve. then youll have your x an y

OpenStudy (anonymous):

im back

OpenStudy (anonymous):

its C?

OpenStudy (anonymous):

no check your answers again

OpenStudy (anonymous):

I got -2, 3

OpenStudy (anonymous):

so D

OpenStudy (anonymous):

yep